OpenStudy (anonymous):

f(x)=definite integral 1/(4+e^t)dt, [0, ln x], with x>0. find (a) f(1) (b) f'(1) (c) inverse of f(0)

OpenStudy (irishboy123):

\[f(x)=\int\limits_{t=0}^{\ln x} dt \qquad \dfrac{1}{4+e^t} \]

jimthompson5910 (jim_thompson5910):

Hint for part (a) \[\Large f(x)=\int\limits_{t=0}^{t=\ln x} \dfrac{1}{4+e^t}dt\] \[\Large f({\color{red}{x}})=\int\limits_{t=0}^{t=\ln ({\color{red}{x}})} \dfrac{1}{4+e^t}dt\] \[\Large f({\color{red}{1}})=\int\limits_{t=0}^{t=\ln ({\color{red}{1}})} \dfrac{1}{4+e^t}dt\]

OpenStudy (anonymous):

so that mean that f(1) = 0

jimthompson5910 (jim_thompson5910):

why is f(1) equal to 0?

jimthompson5910 (jim_thompson5910):

you are correct, but I want to make sure you are using the right integral property

OpenStudy (anonymous):

it is because The Second Fundamental Theorem of Calculus If ƒ is continuous on an open interval I containing a, then for every x in the interval, ƒ(t) dt = ƒ(x)

jimthompson5910 (jim_thompson5910):

no, it's because of this property \[\Large \int_{a}^{a}f(x)dx = 0\]

jimthompson5910 (jim_thompson5910):

the area under the curve from x = a to x = a is going to be 0 because the width of this region is 0 units wide

jimthompson5910 (jim_thompson5910):

think of an infinitely thin strip or line with 0 area

OpenStudy (anonymous):

but it is going from x=o to x= ln(1)

jimthompson5910 (jim_thompson5910):

and ln(1) = 0

jimthompson5910 (jim_thompson5910):

\[\Large \log_b(1) = 0\] for any valid base b

OpenStudy (anonymous):

yes , you're right

jimthompson5910 (jim_thompson5910):

so \[\Large f({\color{black}{1}})=\int\limits_{t=0}^{t=\ln ({\color{black}{1}})} \dfrac{1}{4+e^t}dt\] turns into \[\Large f({\color{black}{1}})=\int\limits_{t=0}^{t=0} \dfrac{1}{4+e^t}dt\]

jimthompson5910 (jim_thompson5910):

for part (b), you'll need to use the fundamental theorem of calculus http://aleph0.clarku.edu/~ma120/FTC.jpg the second part of it. Don't forget to use the chain rule

OpenStudy (anonymous):

for part b, is it also 0

jimthompson5910 (jim_thompson5910):

tell me what you got for \(\Large f \ '(x)\)

jimthompson5910 (jim_thompson5910):

this page may help when it comes to the chain rule: http://www.sosmath.com/calculus/integ/integ03/integ03.html

jimthompson5910 (jim_thompson5910):

Look at the part that has \[\LARGE F(x) = \int_{a}^{u(x)}f(t)dt\]

jimthompson5910 (jim_thompson5910):

notice just under that is \[\LARGE F \ '(x) = f{\Huge(}u(x){\Huge)}u \ '(x)\]

jimthompson5910 (jim_thompson5910):

`for part b, is it also 0` no, the answer to part (b) is NOT zero

OpenStudy (anonymous):

so part b is 1/5

jimthompson5910 (jim_thompson5910):

how did you get 1/5

jimthompson5910 (jim_thompson5910):

what f ' (x) function did you get?